Jen rode her bike 63 miles. She rode twice as many miles on paved road as she did on dirt road. How many miles did she ride on paved road?
Answer by checkley75(3666)

2X+X=63
3X=63
X=63/3
X=21 MILES OF DIRT ROAD
2*21=42 MILES OF PAVED ROADS